OpenStudy (anonymous):

The dimensions of a rectangular prism are given by 5 cm, 6 cm, and 7 cm. Find the change in volume of the prism if the measure of width is doubled. A. The volume is the same as the original volume. B. The volume is 2 times the original volume. C. The volume is 3 times the original volume. D. The volume is 4 times the original volume. I chose letter B

OpenStudy (anonymous):

@AccessDenied

OpenStudy (kainui):

V=l*w*h but if the width doubles, then we replace w with (2w) V=l*2w*h=2(l*w*h) So we can see that it doubles the entire volume!

OpenStudy (anonymous):

so B is right ? @kainui